Q1: Your Company has two alternative opportunities, each requiring your entire capital
investment budget of $300,000. Alternative A will retum $400,000 at the end of one year.
Alternative B will retum $150,000 at the end of each of the first three years. Assuming that your
company requires 15 percent return on investment before taxes. Which (if either) alternative should
you recommend on the basis of (a) simple payback time? (b) Net present worth?




Q2: Briefly describe different phases of Technology Lifecycle. Give examples of products in each phase.


help me please
Answer #1

Q1.

(a) Simple payback time

Simple payback time for alternative A = Investment / annual return

= 300000 / 400000

= 0.75 year

= 9 months

Simple payback time for alternative B = 300000 / 150000

= 2 years

= 24 months

(b) Net Present Worth

Net Present Worth of alternative A = 400000*1/1.15 - 300000

= $ 47,826.09

Net Present Worth of alternative B = 150000 *(1/1.15+1/1.15^2+1/1.15^3) - 300000

= $ 42,483.77

We see that Simple Payback time of alternative A is lower, hence better.

and, Net Present Worth of alternative A is higher, hence better

On the basis of both measures, alternative A is better.

Therefore, we should recommended alternative A